OpenStudy (linn99123):

Need help super fast will fan and give medal Solve. 8y – 24 > –16 A. y > –5 B. y > –1 C. y > 1 D. y > 5

OpenStudy (anonymous):

8y+16=24 -16 from both sides 8y=8 divide both sides by 8 y=1

OpenStudy (anonymous):

Yes, it is.. @jjuden

OpenStudy (campbell_st):

@jjuden... you're answer is incorrect...

OpenStudy (anonymous):

Simplifying 8y + -24 = -16 Reorder the terms: -24 + 8y = -16 Solving -24 + 8y = -16 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '24' to each side of the equation. -24 + 24 + 8y = -16 + 24 Combine like terms: -24 + 24 = 0 0 + 8y = -16 + 24 8y = -16 + 24 Combine like terms: -16 + 24 = 8 8y = 8 Divide each side by '8'. y = 1 Simplifying y = 1

OpenStudy (linn99123):

umm im confuzzled o.o

OpenStudy (campbell_st):

@mikaela.mellish you are required to solve an inequality... not an equation.

OpenStudy (anonymous):

Its c..

OpenStudy (mathmale):

I urge you to show your work, so that others can provide you with meaningful feedback. 8y – 24 > –16 Add 24 to both sides. Linn?

OpenStudy (campbell_st):

now... why give answers... it idea of the site is to assist learning

OpenStudy (mathmale):

Linn?

OpenStudy (linn99123):

hold up im tryign to read everyones response and trying to comprehend it

OpenStudy (anonymous):

It would be c because you'd take out the equals sign and put the greater sign or less than sign in to take its place. :)

OpenStudy (mathmale):

@linn99123 : Please complete this addition of 24 to both sides: 8y – 24 > –16 +24 +24 -------------

OpenStudy (linn99123):

ok i will

OpenStudy (linn99123):

i got 8

OpenStudy (anonymous):

8y-24>=-16 Rearrange : Rearrange the equation by subtracting what is to the right of the greater equal sign from both sides of the inequality : 8*y-24-(-16)≥0 Step by step solution : Step 1 : Pulling out like terms : 1.1 Pull out like factors : 8y - 8 = 8 • (y - 1) Equation at the end of step 1 : 8 • (y - 1) ≥ 0 Step 2 : Solve 8•(y-1) ≥ 0

OpenStudy (anonymous):

I think it may be 8 actually.

OpenStudy (mathmale):

8y – 24 > –16 +24 +24 ------------- 8y > 8 @linn99123 : Please solve for y.